### v1

# Agentic Very Long Video Understanding

Abstract
The advent of always-on personal AI assistants, enabled by all-day wearable devices such as smart glasses, demands a new level of contextual understanding, one that goes beyond short, isolated events to encompass the continuous, longitudinal stream of egocentric video. Achieving this vision requires advances in long-horizon video understanding, where systems must interpret and recall visual and audio information spanning days or even weeks. Existing methods, including large language models and retrieval-augmented generation, are constrained by limited context windows and lack the ability to perform compositional, multi-hop reasoning over very long video streams. In this work, we address these challenges through EGAgent, an enhanced agentic framework centered on entity scene graphs, which represent people, places, objects, and their relationships over time. Our system equips a planning agent with tools for structured search and reasoning over these graphs, as well as hybrid visual and audio search capabilities, enabling detailed, cross-modal, and temporally coherent reasoning. Experiments on the EgoLifeQA and Video-MME (Long) datasets show that our method achieves state-of-the-art performance on EgoLifeQA (57.5%) and competitive performance on Video-MME (Long) (74.1%) for complex longitudinal video understanding tasks.
Aniket Rege at <>, Hyo Jin Kim at <>
1 Introduction
Unlocking always-on personal AI assistants requires understanding not just isolated events, but a continuous stream of evolving user experiences. The recent emergence of AI-equipped wearable consumer devices such as the Ray-Ban Meta glasses, Amazon Echo Frames and Snapchat Spectacles as well as various prototypes (engel2023project; xu2025designing) creates an opportunity for AI agents to maintain persistent access to what users see and do over time. For such assistants to provide helpful, personalized, and context-aware assistance, they need to possess a longitudinal video understanding, i.e. recall and interpret a user’s lived experience over extremely long periods of time (days and months).
In this work, we address the challenge of “very long video understanding”. In prior literature, the definition of “long” has been continuously evolving. Popular benchmarks like MSR-VTT (xu2016msrvtt) and DiDeMo (hendricks2017didemo) where videos are up to a minute in length were once considered long, but recent works have further pushed this frontier to several minutes (Wu2024LongVideoBenchAB; grauman2022ego4d) and up to an hour (fu2025videomme; Zhou2024MLVUBM; wang2025lvbench). The recent EgoLife (yang2025egolife) pushes this frontier to 50 hours of Egocentric video over the course of a week, which is the length we define as very long. Unlike previous benchmarks that focus on large numbers of short, independent videos, EgoLife offers continuous, longitudinal first-person video from six individuals. This week-long horizon enables new research directions, such as tracking entities and their interactions across multiple days, analyzing repeated behaviors and habits, and handling extended periods of inactivity or “lulls” in the video stream. Agentic approaches, which equip agents with tools to search, retrieve, and reason over large corpora, have shown potential in addressing some of these limitations (fan2024videoagent; wang2024videoagent; ma2025drvideo; chu2025graphvideoagent). Existing agentic approaches often struggle to maintain coherent reasoning about entities and their relationships over extended temporal horizons, and have difficulty with fine-grained temporal localization such as tracking repeated actions or habits across days (e.g. “how often did I drink water this week?”). Importantly, there is a need for effective linkage between information from different modalities to support richer and more accurate reasoning.
To address these challenges, we propose EGAgent, an enhanced agentic approach that centers on the extraction and use of an entity scene graph from long videos, where nodes represent people, places, and objects, and edges capture their relationships (e.g. uses, interacts with, mentions, talks to). Each node is annotated with temporal intervals indicating when the relation holds. In our system, we equip a planning agent with the ability to search and reason over this entity graph, as well as utilize a visual search tool (SQL + semantic search hybrid) and an audio transcript search tool. As illustrated in figure˜1, the system uses this graph in combination with audio and visual search to locate all shopping-related taxi rides across multiple days and infer who consistently sits next to the user. By leveraging structured representations like entity graphs, our system preserves complex relationships and supports detailed, compositional reasoning over extended timeframes, overcoming the limitations of existing methods.
We evaluate our EGAgent pipeline on the EgoLifeQA benchmark and demonstrate state-of-the-art performance. Notably, EGAgent surpasses the previous state-of-the-art by 32% and 39.7% on the RelationMap and TaskMaster categories respectively, both of which require multi-hop relational reasoning. Our method also achieves competitive results on the Video-MME (Long) benchmark.
To summarize, our contributions are as follows:
We introduce an entity graph representation (section˜3.2) for long video understanding (section˜3.1), enabling structured, cross-modal reasoning over very long time horizons.
We present an agentic framework (section˜3.3) that queries the entity graph along with visual and audio search tools, exceeding previous state-of-the-art performance on EgoLifeQA by 20.6% (section˜4.3).
We perform a detailed ablation study on entity graph construction and agentic tool usage for very long video understanding on EgoLife (section˜4.5 and Section˜11)
2 Related Work
Long Video Understanding with LLMs. The primary challenge in long-video understanding arises from the limited context window of large language models (LLMs), which restricts the amount of visual information processed at once. To address this, prior work focuses on condensing video inputs before LLM inference (tang2025adaptive; lu2025decafnet; liu2025bolt). Frame selection methods reduce input length by retaining only salient frames while preserving key content (wang2025videotree; buch2025flexible; ye2025tstar), whereas visual token compression techniques distill videos into compact token representations that better fit within context limits (shen2025longvu; shu2025video). These approaches can be query-dependent, selecting frames or tokens based on the input query (liu2025hybrid; hu2025m; man2025adacm; diko2025rewind), or query-independent, producing general summaries irrespective of downstream tasks (yang2025pvc; zhao2025accelerating). Other methods adopt sliding-window or hierarchical summarization strategies to maintain long-range context under fixed token budgets (lu2025vited; zhou2024streaming), or to directly extend the context capacity of LLMs themselves (ding2024longrope; liu2023tcra; jin2024llm).
Video Understanding with Graph-based RAG. Retrieval-augmented generation (RAG) mitigates the context limitations of LLMs by retrieving relevant information from external sources (lewis2020rag; gao2023ragsurvey), which has also been extended to multimodal documents and long-video understanding (yuvis2025rag; faysse2025colpali). Traditional RAG operates over isolated text chunks, often losing relational context. To address this, Graph-based RAG methods such as GraphRAG (edge2024local) and LightRAG (guo2024lightrag) leverages knowledge graphs built from extracted entities and relations from the text corpus. More recently, researchers have begun to explore multi-modal RAG approaches, such as retrieving image frames directly instead of retrieving pre-generated video captions (reddy2025video; wan2025clamr). This approach preserves visual details that may be lost in textual abstraction, enabling more precise and comprehensive responses to complex queries. For instance, Video-RAG (luo2025videorag) performs multi-modal RAG on video frames, automatic speech recognition (ASR) results, optical character recognition (OCR) results, and object-detection results. However, directly retrieving frames also introduces new challenges, including the need for efficient and accurate indexing, retrieval mechanisms, and effective data representations (reddy2025video; wan2025clamr). VideoRAG (ren2025videorag) combines text-, visual-, and graph-based clip retrieval, matching queries to entity descriptions within a graph. AdaVideoRAG (xue2025adavideorag) adaptively selects between no retrieval, naive retrieval, and graph-based retrieval based on question difficulty. RAVU (malik2025ravu) uses VLMs to detect entities, generate frame descriptions, build spatio-temporal graphs, and infer answers. GraphVideoAgent (chu2025graphvideoagent) iteratively retrieves relevant frames via caption-derived graphs. VideoMindPalace (huang2025building) constructs layered spatio-temporal graphs encoding indoor layouts and activity zones, though its reliance on room-level structure limits robustness in open-ended scenes.
Many of these methods either overlook temporal relationships or construct graphs for the entire video at once. In contrast, we introduce an entity graph where each node is annotated with temporal information, making the graph time-aware and allowing it to be incrementally constructed as new data arrives. Experimentally, our method matches the performance of AdaVideoRAG (xue2025adavideorag) on Video-MME (Long) while processing over ten times fewer frames.
Agentic Video Understanding. Recent advances in agentic video understanding have focused on developing systems that can autonomously perceive, reason, and act based on video content (chen2025lvagent). VideoAgent (wang2024videoagent) introduces an agent-based framework where the agent is tasked with iteratively finding the relevant frames in the video for VQA if the information in the initial frames is not sufficient to answer the question. VideoAgent (fan2024videoagent) iteratively employs tools such as object memory search and video-segment search based on video captions and visual embeddings to reach an answer. DrVideo (ma2025drvideo) reframes long-video understanding as long-document understanding by converting videos into text documents, iteratively augmenting them with key frame information and agent-based searches until enough information is gathered for chain-of-thought prediction.
Our proposed EGAgent advances agentic video understanding by integrating a temporally-annotated entity scene graph into the tool-calling loop. Unlike prior systems that rely on unstructured captions or repeated frame retrieval, our approach enables efficient cross-modal search and compositional reasoning for complex, longitudinal queries.
3 Method
Here we formalize the task of very long video understanding (section˜3.1) and extracting entity graph representations of such long videos (section˜3.2). Lastly, we discuss the design of our proposed agentic framework EGAgent which utilizes these entity graph representations for very long video understanding (section˜3.3).
3.1 Task Setup
We focus on the task of Very Long Video Understanding, specifically on video question-answering over videos that potentially span an entire week. Let denote the video sampled at 1 FPS (frame per second). Similarly, let denote the set of transcribed speech with associated time-stamps . At test time, the system receives a complex query in natural language, and must produce a textual answer . Formally, the task is to obtain a mapping .
Naively feeding all frames and transcripts into a multimodal LLM or VLM for such very long videos is infeasible due to context window limitations. The prevailing approach, Video Retrieval Augmented Generation (RAG) (luo2025videorag), first selectively retrieves a small subset of frames and audio transcripts deemed relevant to the user query and conditions the VLM on this retrieved set to generate the answer . However, a naive RAG approach over very long egocentric videos is insufficient to answer egocentric queries which are often entity-centric and require multi-hop reasoning across days. These include tracking repeated behaviors, or interactions between specific people, objects, and locations. Direct embedding based retrieval over unstructured clips or captions struggle to maintain coherent entity identities over time to support compositional constraints such as “all times I talked to person X this week”.
We address this in two steps. First, to support queries over entity relations over time, we construct an entity-centric scene graph that explicitly encodes people, objects, locations, temporally localized relations, and provide a structured index to allow narrowing down to the relevant regions of the video (section˜3.2). Second, we propose an agentic framework EGAgent which involves a planning agent that iteratively decomposes into sub-tasks and invokes specialized retrieval tools including the above constructed entity graph (section˜3.3).
3.2 Entity Graph Representations
From our observations, baseline methods often struggle with questions that require understanding a person’s habits or repeated behaviors over time (e.g., “What do I often check on my phone in the morning?”), as well as those that involve reasoning about interactions and relationships between different entities, such as people, objects, or places across extended periods (e.g., “Before we went to see the dog, who went with me to the second floor to find Tasha?”). Because these methods do not explicitly model entity relationships or track long-term behavioral patterns, their performance on such questions, especially over long time horizons, is limited.
To address this, we construct an entity graph to capture relationships and interactions, enabling our planning agent to query this graph during inference.
Nodes (): entities (i.e., individuals, objects, places)
Edges (): relationships (i.e. interacts with, mentions, talks to, uses), and temporal information
Each edge is annotated with temporal information, allowing us to track the existence, sequence and duration of the corresponding relationships. Such temporal structure is crucial for reasoning about events and interactions that unfold or repeat across long horizons.
Entity Graph Creation. We construct an entity graph from a given collection of text documents which includes audio transcripts, scene descriptions, predicted scene locations (illustrated in figure˜3). We discuss details of extracting scene data to generate these documents in Section˜12. For each document , we apply an LLM-based extractor to jointly identify entities and their relationships:
Here, is the set of entities and is the set of relationships extracted from . The overall entities and relationships are aggregated as:
We assign each node a type to be one of “person”, “object”, “location”. We initially represent each edge as a tuple , where and are the source and target nodes, and is the relationship type. The set of relationship types is:
Each edge is subsequently annotated with temporal information derived from the source document . After temporal annotation, each edge is represented as:
The resulting graph is stored as a set of tuples:
is the supporting text snippet from which the edge was extracted. The graph is stored in memory as a SQLite3 database, with each row corresponding to one tuple. The graph construction process supports incremental updates as new documents arrive, allowing to grow and refine over time.
3.3 Agentic Framework EGAgent
Given the very long video and entity graph representation described above, we propose an agentic framework EGAgent for multi-modal reasoning, summarized in algorithm˜1 and illustrated in figure˜2. EGAgent consists of six main components: a Planning Agent, three Retriever Tools (Visual Search, Audio Transcript Search, and Entity Graph Search), an Analyzer Tool, and a VQA Agent (see ① in figure˜2). We discuss more details of our agent design and provide qualitative examples in section˜9.
Each component operates over a specific data modality or reasoning step. The Planning Agent decomposes a complex user query into sub-tasks, selects appropriate tools, and maintains a working memory that accumulates cross-modal evidence. Retriever Tools (Visual Search, Audio Transcript Search, Entity Graph Search) access different data sources to find relevant information for each sub-task, the Analyzer Tool filters and distills retrieved information, and the VQA Agent produces the final answer from the accumulated evidence.
Planning Agent orchestrates the entire reasoning process. Given a user query along with natural language definitions for each tool, Planning Agent performs a joint decomposition of into a sequence of sub-tasks each sub-task with an associated and with appropriate query arguments (Lines 2-3 in Algorithm˜1).
Each sub-task targets a specific aspect of the information needed such as object localization, checking diarized speech, or confirming past interaction. For each , the Planning Agent selects a retriever tool from one of the following: (i) Visual Search Tool () retrieves visual content. (ii) Audio Transcript Search Tool () retrieves transcribed speech. (iii) Entity Graph Search Tool () queries an entity-centric scene graph. The retrieved content is passed to the Analyzer Tool and the corresponding analysis is updated to the working memory. . Such iterative process allows EGAgent to progressively refine its understanding of the query while keeping per-sub-task context size manageable. Finally, the VQA Agent consumes the working memory and original query to provide a final answer. See ② in figure˜2 for an example of the planning agent reasons over cross-modal information retrieved with retriever tools.
Visual Search Tool samples video frames at 1FPS and embeds each frame as using a vision-encoder (tschannen2025siglip). The generated embeddings along with attributes such as timestamp, location are stored in a vector database which supports efficient retrieval. At inference, the Planning Agent provides a text sub-query (embedded as ) and optional attribute filters (e.g. “kitchen", “morning”). The tool computes cosine similarity for filtered rows in the vector database returning the -nearest neighbors for further analysis.
Audio Transcript Search Tool operates over text transcripts. We consider two variants (i) LLM-based search where we feed entire transcripts to an LLM for a relevant time range (parallelized over days due to context limits) (ii) BM25-based lexical search. The former provides significantly better quality results at the cost of higher latency.
Entity Graph Search Tool queries the entity-centric scene graph introduced in section˜3.2 and stored tuples in a SQLite database (Equation˜5). During inference, the Planning Agent issues SQL queries over the following fields: (i) time filter (ii) keyword text search (iii) entity source and/or target nodes and (iv) relationship type . In practice, real-world data is often incomplete or noisy, so the Planning Agent adopts a “strict-to-relaxed” query strategy: it first issues an exact match query on all specified fields, and if no results are found, incrementally relaxes constraints by broadening the time window, allowing partial text matches, and finally relaxing the relationship type filter. This strategy maximizes precision when possible while increasing recall when exact matches are unavailable (see ③ in figure˜2 for an example query trace and section˜10 for qualitative examples of SQL querying).
Analyzer Tool determines the relevance of the retrieved context for each sub-task via an LLM to perform lightweight reasoning, evidence extraction, and optional de-duplication.
VQA Agent is a multi-modal LLM that conditions on and the compact evidence in to generate the final answer (Algorithm 1, Line 13), enabling detailed, temporally coherent reasoning over week-long egocentric videos.
4 Experiments
We evaluate EGAgent against baselines on two benchmarks, EgoLifeQA and Video-MME (Long), which focus on very long video understanding. Here we discuss implementation details of our EGAgent and analyze its performance on these datasets. Lastly, we discuss a few ablation studies on entity graph extraction and wall-clock latency.
4.1 Evaluation Benchmarks
EgoLifeQA: EgoLifeQA consists of 500 long-context Multiple-Choice Questions (MCQs) derived from the EgoLife (yang2025egolife) dataset, in which six participants lived together for one week, continuously recording their daily activities using Project Aria glasses (engel2023project). The benchmark focuses on the 50 hours of videos taken from the perspective of Jake, one of the six participants. The MCQs cover practical questions such as locating items, recalling past events, tracking habits, and analyzing social interactions. Each question has four candidate answers with a single correct option. Each question is associated with query time (e.g. 11:34 AM on day 4) and a manually verified target time, indicating the specific portion of the video that contains the information needed to answer the MCQ correctly.
Video-MME (Long): Video-MME (fu2025videomme) comprises 900 videos, with 2700 MCQs. The benchmark is divided into Short, Medium, and Long subsets based on video length. We focus on the Long subset that consists of 300 videos that range from 30 to 60 minutes. (section˜4.4).
4.2 Implementation Details
To prepare the entity graphs for our experiments, we extract a separate graph for each video in the Video-MME (fu2025videomme) dataset. For EgoLifeQA (yang2025egolife), due to the increased likelihood of LLM invocation failures with longer input transcripts, we instead extract one graph per hour of video. In both datasets, audio is represented by text transcripts. For Video-MME, transcripts are generated using an ASR foundation model such as Whisper. In contrast, EgoLife provides manually diarized transcripts, which include both speaker identities and the corresponding speech content. We discuss more details and provide code snippets and all agent and tool use prompts in Section˜12.
| Category | Method | # Frames | Modality | MCQ Accuracy (%) | Average Gain (%) | Average # Tokens | |||||
| EL | ER | HI | RM | TM | Average | ||||||
| MLLMs (Uniform Sampling) | LLaVA-Video-7B | 64 | F | – | – | – | – | – | 36.4 | 32K* | |
| GPT-4.1 | 1FPS | C | 32.0 | 39.7 | 39.3 | 32.8 | 39.7 | 36.0 | 285K | ||
| Gemini 2.5 Pro | 3000 | F, T | 45.6 | 48.4 | 51.7 | 41.6 | 52.4 | 46.8 | +9.9 | 807K | |
| RAG | LLaVA-Video-7B + Video-RAG | 64 | F | – | – | – | – | – | 30.0 | 18K* | |
[…]
| EgoButler GPT-4o | 0 | C, T | 34.4 | 42.1 | 29.5 | 30.4 | 44.4 | 36.2 | 19K* | ||
| VideoAgent | 1FPS8 | F | – | – | – | – | – | 29.2 | 128K* | ||
| LLaVA-OneVision-7B + T* | 1FPS8 | F, T | – | – | – | – | – | 35.4 | 32K* | ||
| Ego-R1 Qwen-2.5-3B-Instruct | 1FPS | F, C, T | – | – | – | – | – | 36.0 | 128K* | ||
| Ours | EGAgent GPT-4.1 (F + T) | 1FPS50 | F, T | 48.0 | 48.4 | 55.7 | 40.0 | 61.9 | 48.6 | +11.7 | 551K |
| EGAgent GPT-4.1 (EG + F + T) | 1FPS50 | F, C, T | 44.0 | 49.2 | 55.7 | 53.6 | 66.7 | 50.7 | +13.8 | 571K | |
| EGAgent GPT-4o (EG + F + T) | 1FPS50 | F, C, T | 44.8 | 54.8 | 59.0 | 44.0 | 61.9 | 44.6 | +7.7 | 652K | |
| EGAgent Gemini 2.5 Pro (EG + F + T) | 1FPS50 | F, C, T | 54.4 | 57.1 | 60.3 | 62.4 | 74.6 | 57.5 | +20.6 | 880K | |
4.3 Analysis on EgoLifeQA Benchmark
We compare our approach against various strong baselines in three categories: 1) MLLM with uniform sampling; 2) MLLM with RAG; and 3) existing agentic approaches.
Baselines. To handle extremely long videos in EgoLifeQA, frame sampling in MLLM baselines varies based on their respective context window size. GPT-4.1 takes video captions that were generated for every 30-second video snippet sampled at 1 FPS. We sample 3000 frames uniformly along with the audio transcripts for Gemini 2.5 Pro. The results of LLaVa-Video-7B (zhang2025llavavideo) and LLaVA-Video-7B combined with Video-RAG (luo2025videorag) are reported in yang2025egolife.
We compare our approach with the following existing agentic methods: EgoButler (yang2025egolife), a hierarchical text-based Retrieval-Augmented Generation (RAG) approach, and Ego-R1 (tian2025egor1), a lightweight 3B-parameter agent trained on egocentric data, including portions of EgoLife for tool calling. We report results of all RAG and existing agentic approaches in table˜1 directly from these works.
Performance Analysis. table˜1 presents a comprehensive comparison of methods on the EgoLifeQA benchmark. Our proposed agentic system, which incorporates entity graph reasoning, achieves strong performance across all evaluation categories and establishes a new state-of-the-art. Notably, while Gemini 2.5 Pro with uniform sampling already outperforms the previous best results (EgoButler), our agentic system based on Gemini 2.5 Pro delivers an additional improvement of 10.7%, highlighting the significant value of entity graph reasoning.
Furthermore, the benefits of entity graph reasoning are not limited to a single MLLM backbone. Applying the same agentic framework to the GPT-4.1 backbone also yields notable gains over its uniform sampling counterpart. These results demonstrate that integrating entity graph reasoning within agentic systems consistently enhances performance on very long video understanding tasks.
To compare with existing agentic systems, we run our agentic system on the same LLM backbone (GPT-4o) with other agentic system. Our entity graph agent consistently surpasses other agentic approaches utilizing the same model, including EgoButler (+8.4%), VideoAgent (+15.4%), and Ego-R1 (+8.6%).
table˜1 demonstrates that, when using the same backbone, the proposed method incorporating an entity graph substantially improves performance on EgoLifeQA. It outperforms the baseline (without entity graph) in 4 out of 5 categories, with particularly notable gains in the RelationMap and TaskMaster categories. This improvement can be attributed to the entity graph’s ability to enable cross-modal reasoning.
figure˜4 further illustrates the performance gap between the proposed agentic approach and Gemini 2.5 Pro with uniform sampling. It is evident that the agentic system benefits the most on RelationMap and TaskMaster categories, which require multi-hop relational reasoning. Specifically, our approach surpasses the previous state-of-the-art and Gemini 2.5 Pro by 32% and 20.8%, respectively, on RelationMap QAs, and achieves impressive gains of 39.7% and 17.5% in the TaskMaster category. We discuss more examples and benchmark analyses in section˜11.
| Category | Method | Context | # Frames | Modality | Accuracy (%) | # Tokens | |
| MLLMs (Uniform Sampling) | Gemini 2.5 Pro | 1M | 256 | F, A | 82.0 | 100K* | |
| GPT-4.1 | 1M | 384 | F | 72.0 | 60K* | ||
| RAG | Video-RAG (Qwen2.5-VL-7B) | 32K | 32 | F, O, T | 43.3 | 10K* | |
| AdaVideoRAG (Qwen2.5-VL-7B) | 128K | 768 | F, C, T | 47.7 | 128K* | ||
| Agentic Baselines | DrVideo (DeepSeek V2.5) | 128K | 0.2 FPS | F, T | 71.7 | 128K* | |
32K | 32 | F, T | 72.4 | 32K* | ||
| Ours | EGAgent (Qwen2.5-VL-7B) | 32K | 1FPS50 | F, C, T | 47.8 | 172K | |
[…]
| Method | VLM | EGX | # F | T Search | Accuracy (%) | Gain (%) |
| EgoButler | GPT-4o | – | 0 | LLM | 36.2 | - |
| EGAgent | GPT-4o | T | 50 | BM25 | 36.6 | +0.4 |
| T+C | 50 | BM25 | 39.4 | +3.2 | ||
| T+C | 50 | LLM | 44.6 | +8.4 | ||
| GPT-4.1 | T | 0 | - | 36.8 | +0.6 | |
| T | 50 | BM25 | 42.2 | +6.0 | ||
| T | 50 | LLM | 49.2 | +13.0 | ||
| T+C | 50 | BM25 | 43.9 | +7.7 | ||
| T+C | 50 | LLM | 50.7 | +14.5 | ||
| Gemini 2.5 Pro | T | 50 | BM25 | 48.6 | +12.4 | |
| T+C | 50 | BM25 | 51.8 | +15.6 | ||
| T+C | 50 | LLM | 57.5 | +21.3 |
4.4 Analysis on Video-MME Benchmark
We also evaluate our entity graph agent on the long subset of Video-MME in table˜2. Because Gemini 2.5 Pro can process native video (frames + audio) without the need for transcripts, it remains the state-of-the-art in this sub-hour length regime. Using an identical LLM backbone (Qwen2.5-VL-7B), EGAgent surpasses Video-RAG (+4.5%), and matches the performance of AdaVideoRAG while processing over fewer frames.
Compared with recent agentic approaches (guo2025deepseek; yuan2025videodeepresearch) that use frontier models as their LLM backbone, our EGAgent with a Gemini 2.5 Pro backbone demonstrates strong performance, second only to native Gemini 2.5 Pro that processes 256 frames. In contrast, EGAgent uses only a fifth of the image frames compared to the baseline. More importantly, uniformly sampling with MLLMs like Gemini 2.5 Pro does not scale well to extremely long videos, as demonstrated in the EgoLifeQA benchmark section˜4.3.
| Method | T Search | Accuracy (%) | Runtime (sec) | #Tokens |
| EGAgent GPT-4.1 | BM25 | 43.9 | 125 | 172K |
| LLM | 50.7 | 169 | 571K |
4.5 Ablations
Extraction of Entity Graph. We compare two variants of Entity Graph Extraction (EGX) in EgoLifeQA in table˜3. The additional information from visual captions increases MCQ accuracy by 2.6% on average across all three MLLM backbones (GPT-4o, GPT-4.1 and Gemini 2.5 Pro).
Agent Wall-Clock Latency. We tabulate the wall-clock latency of our EGAgent pipeline in table˜4. EGAgent takes between two and three minutes to answer an MCQ, depending on the number of sub-tasks required by the planning agent. We also evaluate the latency impact of the transcript search and replace our default LLM search with BM25 (robertson2009bm25), which drops token usage by at the cost of a 6.8% MCQ accuracy drop on average.
We discuss more ablations on tool usage and retrieval quality in Section˜11.
5 Conclusion
We introduce a novel EGAgent framework (section˜3.3) for longitudinal video understanding, addressing the unique challenges posed by always-on personal AI assistants processing very long egocentric video streams. By leveraging entity scene graphs (section˜3.2) and specialized tools for structured, cross-modal reasoning, our approach enables detailed and temporally coherent analysis. Experiments on EgoLifeQA (section˜4.3) and Video-MME (Long) (section˜4.4) demonstrate state-of-the-art performance on tasks requiring the tracking of entities, behaviors, and relationships over extended periods. As video lengths continue to grow, we believe our results highlight the potential of agentic planning over structured representations of inter-entity relationships for very long video understanding moving forward.
6 Limitations
While our EGAgent achieves strong performance on longitudinal video understanding tasks, it is important to note that the construction of entity scene graphs depends on the accuracy of upstream perception and language models, which may occasionally introduce errors in extracting entities and relationships. Additionally, our experiments relied on transcripts and for EgoLife, manually annotated speaker diarization. In scenarios where off-the-shelf diarization models are used, downstream performance is likely to be adversely affected by prediction errors.
7 Ethical Considerations
Our work uses the publicly available EgoLife dataset, which was released under an MIT license. We adhere to all terms of use associated with this dataset. The EgoLife dataset automatically detects and blurs faces and other personally identifiable information (PII) such as sensitive audio content. We also use the Video-MME dataset, which was released under a custom license111License: https://github.com/MME-Benchmarks/Video-MME. We have adhered to all terms of use associated with this dataset, using an unmodified version strictly for academic research. In addition to these pre-existing safeguards, we have taken extra care to protect individual privacy in our reporting: all faces appearing in the figures throughout this paper have been manually blurred.

### v2

@@
 Experiments on the EgoLifeQA and Video-MME (Long) datasets show that our method achieves state-of-the-art performance on EgoLifeQA (57.5%) and competitive performance on Video-MME (Long) (74.1%) for complex longitudinal video understanding tasks.
-Aniket Rege at <>, Hyo Jin Kim at <>
+[…]
 1 Introduction
@@
 The recent emergence of AI-equipped wearable consumer devices such as the Ray-Ban Meta glasses, Amazon Echo Frames and Snapchat Spectacles as well as various prototypes (engel2023project; xu2025designing) creates an opportunity for AI agents to maintain persistent access to what users see and do over time.
-For such assistants to provide helpful, personalized, and context-aware assistance, they need to possess a longitudinal video understanding, i.e.
-recall and interpret a user’s lived experience over extremely long periods of time (days and months).
+For such assistants to provide helpful, personalized, and context-aware assistance, they need to possess longitudinal video understanding, i.e., the ability to recall and interpret a user’s lived experience over extremely long periods of time (days and months).
 In this work, we address the challenge of “very long video understanding”.
@@
 Popular benchmarks like MSR-VTT (xu2016msrvtt) and DiDeMo (hendricks2017didemo) where videos are up to a minute in length were once considered long, but recent works have further pushed this frontier to several minutes (Wu2024LongVideoBenchAB; grauman2022ego4d) and up to an hour (fu2025videomme; Zhou2024MLVUBM; wang2025lvbench).
-The recent EgoLife (yang2025egolife) pushes this frontier to 50 hours of Egocentric video over the course of a week, which is the length we define as very long.
+The recent EgoLife (yang2025egolife) pushes this frontier to beyond 50 hours of Egocentric video over the course of a week, which is the length we define as very long.
 Unlike previous benchmarks that focus on large numbers of short, independent videos, EgoLife offers continuous, longitudinal first-person video from six individuals.
@@
 Agentic approaches, which equip agents with tools to search, retrieve, and reason over large corpora, have shown potential in addressing some of these limitations (fan2024videoagent; wang2024videoagent; ma2025drvideo; chu2025graphvideoagent).
-Existing agentic approaches often struggle to maintain coherent reasoning about entities and their relationships over extended temporal horizons, and have difficulty with fine-grained temporal localization such as tracking repeated actions or habits across days (e.g.
-“how often did I drink water this week?”).
-Importantly, there is a need for effective linkage between information from different modalities to support richer and more accurate reasoning.
-To address these challenges, we propose EGAgent, an enhanced agentic approach that centers on the extraction and use of an entity scene graph from long videos, where nodes represent people, places, and objects, and edges capture their relationships (e.g.
-uses, interacts with, mentions, talks to).
-Each node is annotated with temporal intervals indicating when the relation holds.
-In our system, we equip a planning agent with the ability to search and reason over this entity graph, as well as utilize a visual search tool (SQL + semantic search hybrid) and an audio transcript search tool.
+Existing agentic approaches often struggle to maintain coherent reasoning about entities and their relationships over extended temporal horizons, and have difficulty with fine-grained temporal localization such as tracking repeated actions or habits across days (e.g., “how often did I drink water this week?”).
+Importantly, effective linkage between information from different modalities is needed to support richer and more accurate reasoning.
+To address these challenges, we propose EGAgent, an enhanced agentic approach that centers on the extraction and use of an entity scene graph from long videos, where nodes represent people, places, and objects, and edges capture their relationships (e.g., uses, interacts with, mentions, talks to).
+Each edge is annotated with temporal intervals indicating when the relation holds.
+In our proposed EGAgent system, we equip a planning agent with the ability to search and reason over this entity graph, as well as utilize a visual search tool (SQL + semantic search hybrid) and an audio transcript search tool.
 As illustrated in figure˜1, the system uses this graph in combination with audio and visual search to locate all shopping-related taxi rides across multiple days and infer who consistently sits next to the user.
-By leveraging structured representations like entity graphs, our system preserves complex relationships and supports detailed, compositional reasoning over extended timeframes, overcoming the limitations of existing methods.
+By leveraging structured representations like entity graphs, EGAgent preserves complex relationships and supports detailed, compositional reasoning over extended timeframes, overcoming the limitations of existing methods.
 We evaluate our EGAgent pipeline on the EgoLifeQA benchmark and demonstrate state-of-the-art performance.
 Notably, EGAgent surpasses the previous state-of-the-art by 32% and 39.7% on the RelationMap and TaskMaster categories respectively, both of which require multi-hop relational reasoning.
-Our method also achieves competitive results on the Video-MME (Long) benchmark.
+EGAgent also achieves competitive results on the Video-MME (Long) benchmark.
 To summarize, our contributions are as follows:
@@
 Traditional RAG operates over isolated text chunks, often losing relational context.
-To address this, Graph-based RAG methods such as GraphRAG (edge2024local) and LightRAG (guo2024lightrag) leverages knowledge graphs built from extracted entities and relations from the text corpus.
+To address this, Graph-based RAG methods such as GraphRAG (edge2024local) and LightRAG (guo2024lightrag) leverage knowledge graphs built from extracted entities and relations from the text corpus.
 More recently, researchers have begun to explore multi-modal RAG approaches, such as retrieving image frames directly instead of retrieving pre-generated video captions (reddy2025video; wan2025clamr).
@@
 AdaVideoRAG (xue2025adavideorag) adaptively selects between no retrieval, naive retrieval, and graph-based retrieval based on question difficulty.
-RAVU (malik2025ravu) uses VLMs to detect entities, generate frame descriptions, build spatio-temporal graphs, and infer answers.
+RAVU (malik2026ravu) uses VLMs to detect entities, generate frame descriptions, build spatio-temporal graphs, and infer answers.
 GraphVideoAgent (chu2025graphvideoagent) iteratively retrieves relevant frames via caption-derived graphs.
@@
 DrVideo (ma2025drvideo) reframes long-video understanding as long-document understanding by converting videos into text documents, iteratively augmenting them with key frame information and agent-based searches until enough information is gathered for chain-of-thought prediction.
+Similarly SiLVR (zhang2026silvr) operates in the text domain by compressing dense visual captions and using a downstream reasoning LLM for video understanding.
 Our proposed EGAgent advances agentic video understanding by integrating a temporally-annotated entity scene graph into the tool-calling loop.
@@
 Here we formalize the task of very long video understanding (section˜3.1) and extracting entity graph representations of such long videos (section˜3.2).
-Lastly, we discuss the design of our proposed agentic framework EGAgent which utilizes these entity graph representations for very long video understanding (section˜3.3).
+Lastly, we discuss the design of the proposed agentic framework EGAgent which utilizes these entity graph representations for very long video understanding (section˜3.3).
 3.1 Task Setup
@@
 These include tracking repeated behaviors, or interactions between specific people, objects, and locations.
-Direct embedding based retrieval over unstructured clips or captions struggle to maintain coherent entity identities over time to support compositional constraints such as “all times I talked to person X this week”.
+Direct embedding-based retrieval over unstructured clips or captions struggles to maintain coherent entity identities over time to support compositional constraints such as “all times I talked to person X this week”.
 We address this in two steps.
@@
 Because these methods do not explicitly model entity relationships or track long-term behavioral patterns, their performance on such questions, especially over long time horizons, is limited.
-To address this, we construct an entity graph to capture relationships and interactions, enabling our planning agent to query this graph during inference.
+To address this, we construct an entity graph to capture relationships and interactions, enabling the planning agent to query this graph during inference.
 Nodes (): entities (i.e., individuals, objects, places)
-Edges (): relationships (i.e. interacts with, mentions, talks to, uses), and temporal information
+Edges (): relationships (i.e., interacts with, mentions, talks to, uses), and temporal information
 Each edge is annotated with temporal information, allowing us to track the existence, sequence and duration of the corresponding relationships.
@@
 The generated embeddings along with attributes such as timestamp, location are stored in a vector database which supports efficient retrieval.
-At inference, the Planning Agent provides a text sub-query (embedded as ) and optional attribute filters (e.g.
-“kitchen", “morning”).
+At inference, the Planning Agent provides a text sub-query (embedded as ) and optional attribute filters (e.g., “kitchen", “morning”).
 The tool computes cosine similarity for filtered rows in the vector database returning the -nearest neighbors for further analysis.
@@
 We evaluate EGAgent against baselines on two benchmarks, EgoLifeQA and Video-MME (Long), which focus on very long video understanding.
-Here we discuss implementation details of our EGAgent and analyze its performance on these datasets.
+Here we discuss implementation details of EGAgent and analyze its performance on these datasets.
 Lastly, we discuss a few ablation studies on entity graph extraction and wall-clock latency.
@@
 Each question has four candidate answers with a single correct option.
-Each question is associated with query time (e.g.
-11:34 AM on day 4) and a manually verified target time, indicating the specific portion of the video that contains the information needed to answer the MCQ correctly.
+Each question is associated with query time (e.g., 11:34 AM on day 4) and a manually verified target time, indicating the specific portion of the video that contains the information needed to answer the MCQ correctly.
 Video-MME (Long): Video-MME (fu2025videomme) comprises 900 videos, with 2700 MCQs.
@@
 4.3 Analysis on EgoLifeQA Benchmark
-We compare our approach against various strong baselines in three categories: 1) MLLM with uniform sampling; 2) MLLM with RAG; and 3) existing agentic approaches.
+We compare the proposed EGAgent against strong baselines in three categories: 1) MLLM with uniform sampling; 2) MLLM with RAG; and 3) existing agentic approaches.
 Baselines.
@@
 The results of LLaVa-Video-7B (zhang2025llavavideo) and LLaVA-Video-7B combined with Video-RAG (luo2025videorag) are reported in yang2025egolife.
-We compare our approach with the following existing agentic methods: EgoButler (yang2025egolife), a hierarchical text-based Retrieval-Augmented Generation (RAG) approach, and Ego-R1 (tian2025egor1), a lightweight 3B-parameter agent trained on egocentric data, including portions of EgoLife for tool calling.
+We compare our EGAgent with the following existing agentic methods: EgoButler (yang2025egolife), a hierarchical text-based Retrieval-Augmented Generation (RAG) approach, and Ego-R1 (tian2025egor1), a lightweight 3B-parameter agent trained on egocentric data, including portions of EgoLife for tool calling.
 We report results of all RAG and existing agentic approaches in table˜1 directly from these works.
@@
 Our proposed agentic system, which incorporates entity graph reasoning, achieves strong performance across all evaluation categories and establishes a new state-of-the-art.
-Notably, while Gemini 2.5 Pro with uniform sampling already outperforms the previous best results (EgoButler), our agentic system based on Gemini 2.5 Pro delivers an additional improvement of 10.7%, highlighting the significant value of entity graph reasoning.
+Notably, while Gemini 2.5 Pro with uniform sampling already outperforms the previous best results (EgoButler), our EGAgent with a Gemini 2.5 Pro backbone delivers an additional improvement of 10.7%, highlighting the significant value of entity graph reasoning.
 Furthermore, the benefits of entity graph reasoning are not limited to a single MLLM backbone.
@@
 These results demonstrate that integrating entity graph reasoning within agentic systems consistently enhances performance on very long video understanding tasks.
-To compare with existing agentic systems, we run our agentic system on the same LLM backbone (GPT-4o) with other agentic system.
+To compare with existing agentic systems, we run EGAgent on the same LLM backbone (GPT-4o) with other agentic system.
 Our entity graph agent consistently surpasses other agentic approaches utilizing the same model, including EgoButler (+8.4%), VideoAgent (+15.4%), and Ego-R1 (+8.6%).
@@
 Agent Wall-Clock Latency.
-We tabulate the wall-clock latency of our EGAgent pipeline in table˜4.
+We tabulate the wall-clock latency of EGAgent pipeline in table˜4.
 EGAgent takes between two and three minutes to answer an MCQ, depending on the number of sub-tasks required by the planning agent.
-We also evaluate the latency impact of the transcript search and replace our default LLM search with BM25 (robertson2009bm25), which drops token usage by at the cost of a 6.8% MCQ accuracy drop on average.
+We also evaluate the latency impact of the transcript search and replace the default LLM search with BM25 (robertson2009bm25), which drops token usage by at the cost of a 6.8% MCQ accuracy drop on average.
 We discuss more ablations on tool usage and retrieval quality in Section˜11.